OpenStudy (anonymous):

Joe Company in 2006 had sales of $800,000. In 2007 sales were up 75 percent. Calculate the sales for 2007

OpenStudy (anonymous):

800,000 * 0.75 + 800,000

OpenStudy (anonymous):

= 600,000 + 800,000 = 1,400,000

OpenStudy (anonymous):

how did you come up with that?

OpenStudy (anonymous):

75% = 75/100 = 0.75

OpenStudy (anonymous):

this is the extra the copany sold in 2007 - on top of 800,000 made in 2006